OpenStudy (anonymous):

x/9=4; Is 14 a solution?

OpenStudy (jamesj):

If x = 14 were a solution to x/9 = 4 then it must be the case that 14/9 =4. Is that true?

OpenStudy (anonymous):

I think so. This is what I'm guessing.

OpenStudy (jamesj):

What? Are you saying 14/9 = 4?

OpenStudy (anonymous):

divide 14/9 and see what u get?

OpenStudy (anonymous):

I got 1.5

myininaya (myininaya):

if you have x/4=6 and you know 4 * 6=24 then what is x?

OpenStudy (jamesj):

and is that equal to 4? No. It is not the case that 14/9 = 4. Therefore x = 14 is not a solution to x/9 = 4

OpenStudy (anonymous):

24

OpenStudy (anonymous):

Ok I see understand it now.

myininaya (myininaya):

ok so if you want to solve x/9=4 then all you have to do is figure out what blank times blank is

OpenStudy (anonymous):

X/9=4 4x9= 36 X=36